Problem A.1


The graph below is made of three line segments:


-1 1 2 3 4 5 6 7 8 9 10 11 12


1


2


3


4


y


x


f(x)


g(x)


h(x)


The segments correspond to the following three functions:


f(x) = x − 2, g(x) = p


4 − (x − 6)2 + 2, h(x) = x − 6


Find the total length L of the graph between x = 2 and x = 10.

1
Expert's answer
2021-11-24T05:43:07-0500

(I) length of f(x)

Given

"f(x)=x-2, for \ 2\leq x\leq4"


the length would be the length of hypotenuse and right angled isosceles triangle of side length = 2 units

"\implies h^2=b^2+h^2\\\implies h^2=2^2+2^2\\\implies h^2=8\\h=2\sqrt{2}"


length of graph contributed by

"f(x)=2\sqrt{2}"



(ii) Length of g(x)

Given

"g(x)=2\sqrt{4-(x-6)^2+2}, \ for \ 4\leq x \leq8"


"\implies g(x)-2=\sqrt{4-(x-6)^2}\\\implies(g(x)-2)^2=4-(x-6)^2\\\implies(x-6)^2+(g(x)-2)^2=2^2"


so, g(x) is a circle with center at (6,2) and radius = 2 units

length contributed by g(x) would be circumference of semicircle


length contributed by g(x)="\pi r=2\pi"


(III) Length of h(x)

Given,

"f(x)=x-6, for\ 8\leq x\leq10"


the length would be the length of hypotenuse of right angled triangle of base =height=2 units


"\implies h^2=b^2+h^2\\\implies h^2=2^2+2^2\\\implies h^2=8\\h=2\sqrt{2}"


length of graph contributed by

h(x)="2\sqrt{2}"

Total length of the graph

"L=length\ contributed \ by \ f(x)+length\ contributed \ by \ g(x)+length\ contributed \ by \ h(x)" "\implies L=2\sqrt{2}+2\pi+2\sqrt{2}\\\implies L=4\sqrt{2}+2\pi"